Document Page: First | Prev | Next | All | Image | This Release | Search

File: 970107_jul96_decls4_0001.txt
Page: 0001
Total Pages: 3

Subject: CG UPDATE                                                       

Unit: 3D AD       

Parent Organization: VII  CORPS  

Box  ID: BX001340

Folder Title: DAILY STAFF JOURNAL-TOC LOGS-23-24 FEB 91                                                       

Document Number:         30

Folder Seq  #:         34





                                                  IFIED



    *MEMO FOR S3 (MAIN AND TA




    SUBJECT: CG Update 2307onr

            (U) SWO:
            @u la. Hi Pressure dominating the area.
            U@b. All air green for next 72 hours.
            C. Illum tonite - 61%; 5-6 hours of NVG use, until about 0249

    24U)G2:

            a. No change 4n PIR/eneny projection
            t. No significant enemy activity; some repositioning south of
    B.1     d-appear to be concerned about extreme west +lank.
               Some last minute repositioning and further preparation +or chem
    we.rr=,e observed.
            (U)d. DIA reports only 72 days of supply in theatre; after hostilities
    bl,gin, only 20 days of supply with 16 days of CL V.
            (U)e. Lar e mine+ield appeared last niqht--150k mines via PU6045
            u     I
    **C@       ENT: Any mine-fielcis between 2ACR & IS ABC? A;
               but EPW reports some.
            (1119. IEW is incapable of rapid targeting.
            @Vlh. Probable course of action: remain in place; defend in depth.           At
            until they can define the major attack.
            U i. Major areas for collection today: South of Taw; 26th Div
    x.      M.@d COMMENT: Ask questions about mines and imagery in western sector        and
    fr-om LD Melon - Grape --that whole sector Lip t,iere.

            GS:
             Schwarzkop+ visit cancelled
             GENS Waller & Arnold are moving to KKMC today (forward GP)
             JCEOI will freeze at H-12; all freqs and call signs.
             XVIII ABC Activity:
                - moved FSCL to PL Charger
                - insertion of LRS (101) to PL Brown, Grey ind Red postponed until
    41,oday
    *4CS COMMENT: Sure about "hat? Isn't that an irretrievable action?
                - 24th has an AH64 recon to PL Charger.
                - 3ACR had an ARTY Raid last nitL---took out some bldgs, oil tanks,
    but no enemy personnel or equipment.
heir plans in the eastern sector are
    ,:oday... when you find out, track me down.
            (U)e. 2ACR- a UH60 left Dhahran yesterday, but has not been heard from.
    Sendinq out a search party.
            (U)f . _1AD - NSTR
            (U)g. 3AD - AH64 raid still on hold (not cancelled)(***LNO note: This is a
    change -From last night when COL Hawk told me it was cancelled)
    ,h(u)h. IID - All aircraft South of LD; hit 1 or 2 armored vehicles, 1
            L-elL-d vehicle--encounterc?d less enemy thin previously thought. Received
    ii-iital jamming, but worked around it.
    ,,Iyiki. 1CAV - Patrol out at 1930, engaged MTI w/ 155... no BDA.
            j. ICUK)AD - NSTR
            . MOB TAC deploying today--!ID(M) sector.
            COMMENT: TAC itsel@ is jumping forward to be in place for the 2ACR
    attack.
                                       UNCLASSIFIED
            F=qr:.                  -0

Document Page: First | Prev | Next | All | Image | This Release | Search


Document 3 f:/Week-28/BX001340/DAILY STAFF JOURNAL-TOC LOGS-23-24 FEB 91/cg update:01029718422224
Control Fields 17
File Room = jul96_declassified
File Cabinet = Week-28
Box ID = BX001340
Unit = 3D AD
Parent Organization = VII CORPS
Folder Title = DAILY STAFF JOURNAL-TOC LOGS-23-24 FEB 91
Folder Seq # = 34
Subject = CG UPDATE
Document Seq # = 30
Document Date =
Scan Date =
Queued for Declassification = 01-JAN-1980
Short Term Referral = 01-JAN-1980
Long Term Referral = 01-JAN-1980
Permanent Referral = 01-JAN-1980
Non-Health Related Document = 01-JAN-1980
Declassified = 02-JAN-1997